OpenStudy (anonymous):

Solve the system by the elimination method. -2x + y + 6 = 0 2x + y - 8 = 0 When you eliminate x, what is the resulting equation?

OpenStudy (swissgirl):

Ok elimination is kinda like getting rid of one variable. So lets get rid of the x. All we gotta do is add the 2 equations \(-2x+y=-6\) \(2x+y=8\) ok so lets subtract: \( -2x+y=-6\) \( 2x+y=8\) -------------- \( 0x+ 2y=2\) So the resulting equation is 2y=2

OpenStudy (swissgirl):

This group is not so active i bet your questions would get more attn in the math section

OpenStudy (swissgirl):

hmmmm can u post ur question here it will be easier to solve